A solution is made using 30.3 percent by mass CH₂Cl₂ in CHCl₃. At 30 °C, the vapor pressure of pure CH₂Cl₂ is 490 mm Hg, and the vapor pressure of pure CHCl₃ is 260 mm Hg. The normal boiling point of CHCl₃ is 61.7 °C.

A.) What is the mole fraction of CH₂Cl₂ in the solution?

B.) What is the vapor pressure of the solution?

C.) What is the molality of CH2CL2 in the solution?

D.) What is the boiling point of the solution? (Kb= 3.67 celsius /m)
Answer #1

Ans :-

A). Let mass of CHCl3 i.e. solvent = 100 g

So,

Mass of CH2Cl2 = 30.3 g

No. of moles of CHCl3 (nCHCl3) = Given mass of CHCl3 / Gram molar mass of CHCl3 = 100 g / 119.38 g/mol = 0.83766 mol

and

No.of moles of CH2Cl2 (nCH2Cl2) = Given mass of CH2Cl2​​​​​​​ / Gram molar mass of CH2Cl2 = 30.3 g/84.93 g/mol = 0.35676 mol

Therefore,

Mole fraction of CH2Cl2 = nCH2Cl2/(nCHCl3 + nCH2Cl2)

= 0.35676 mol / (0.35676 mol + 0.83766 mol)

= 0.35676 mol / (1.19442 mol)

= 0.30

Therefore, Mole fraction of CH2Cl2 = 0.30

--------------------------------------------------------------------

B). Mole fraction of CHCl3 = 1 - 0.30 = 0.70

Partial vapor pressure of CHCl3 (PCHCl3) = P0CHCl3 x XCHCl3

= 260 mm Hg x 0.70

= 182 mm Hg

Similarly,

Partial vapor pressure of CH2Cl2 (PCH2Cl2) = P0CH2Cl2 x XCH2Cl2

= 490 mm Hg x 0.30

= 147 mm Hg

So, vapor pressure of the solution = PCHCl3 + PCH2Cl2 = 182 mm Hg + 147 mm Hg = 329 mm Hg

Therefore, vapor pressure of the solution = 329 mm Hg

----------------------------------------------------

C). Molality of CH2Cl2 (m) = No. of moles of CH2Cl2​​​​​​​/ Mass of solvent CHCl3 in kg

= 0.35676 mol / 0.100 Kg

= 3.5676 mol/kg

Therefore, Molality of CH2Cl2 solution = 3.57 m

-----------------------------------------------------

D). Tb - T0b = kb.m

Tb - 61.7 °C = (3.67 °C/m).(3.57 m)

Tb - 61.7 °C = (3.67 °C/m).(3.57 m)

Tb - 61.7 °C = 13.10 °C

Tb  = 13.10 °C + 61.7 °C

Tb  = 74.8 °C

Therefore, boiling point of the solution = 74.8 °C